# Measuring eccentricity and gas-induced perturbation from gravitational waves of LISA massive black hole binaries

### v1

@@ Abstract
We assess the possibility of detecting both eccentricity and gas effects (migration and accretion) in the gravitational wave (GW) signal from LISA massive black hole binaries (MBHBs) at redshift $z=1$. Gas induces a phase correction to the GW signal with an effective amplitude ($C_{\rm g}$) and a semi-major axis dependence (assumed to follow a power-law with slope $n_{\rm g}$). We use a complete model of the LISA response, and employ a gas-corrected post-Newtonian in-spiral-only waveform model TaylorF2Ecc. By using the Fisher formalism and Bayesian inference, we explore LISA’s ability to constrain $C_{\rm g}$ together with the initial eccentricity $e_{0}$, the total redshifted mass $M_{z}$, the primary-to-secondary mass ratio $q$, the dimensionless spins $\chi_{1,2}$ of both component BHs, and the time of coalescence $t_{c}$. We find that simultaneously constraining $C_{\rm g}$ and $e_{0}$ leads to worse constraints on both parameters with respect to when considered individually. Assuming a standard thin viscous accretion disc, for $M_{z}=10^{6}~{}{\rm M}_{\sun}$, $q=8$, $\chi_{1,2}=0.9$, and $t_{c}=4$ years, we can confidently measure (with a relative error of $<50$ per cent) an Eddington ratio as small as ${\rm f}_{\rm Edd}\sim 0.1$ for a circular binary while for an eccentric system only ${\rm f}_{\rm Edd}\gtrsim 1$ can be inferred. The minimum measurable eccentricity is $e_{0}\gtrsim 10^{-2.75}$ in vacuum and $e_{0}\gtrsim 10^{-2}$ in the presence of a circumbinary disc. A weak environmental perturbation (${\rm f}_{\rm Edd}\lesssim 1$) to a circular binary can be mimicked by an orbital eccentricity during in-spiral, implying that an electromagnetic counterpart would be required to confirm the presence of an accretion disc.
@@
pubyear: 2024pagerange: Measuring eccentricity and gas-induced perturbation from gravitational waves of LISA massive black hole binariesE
@@ 1 Introduction
The prospect of the observation of gravitational waves (GWs) in the mHz band in the 2030s looks promising following the adoption by ESA of the Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al. 2017; Barack et al. 2019) and with other projects, such as TianQin (Wang et al., 2019) and Taiji (Gong et al., 2021), being developed. One of the primary expected extragalactic sources for LISA are massive black hole binaries (MBHBs) with primary-to-secondary mass ratios $q\lesssim 10$ and total masses between $10^{4}~{}{\rm M}_{\sun}$ and $10^{8}~{}{\rm M}_{\sun}$, which LISA will be able to observe up to redshift $z\sim 20$ (Amaro-Seoane et al., 2017). Another expected source are intermediate/extreme mass ratio inspirals (I/EMRIs; Babak et al. 2017; Amaro-Seoane 2018) with $q\gtrsim 10^{2}$, which can be observed up to $z\lesssim 2$. MBHBs, with their high signal-to-noise ratios (SNRs; Amaro-Seoane et al. 2017), provide exciting opportunities to not only measure source properties with high accuracy but also place constraints on the properties of their environments.
@@
GWs can be an important tool to not only measure the source properties but also probe imprints of the environment in which the binary is evolving. In the coalescence phase, gas mainly affects the binary via migration torques and mass accretion. The detectability of the imprint on the emitted waveform of these effects (see, e.g. Barausse et al., 2014; Garg et al., 2022; Caputo et al., 2020) is dependent on the details of the gas inflow. While most of the works on environmental measurements from GWs have focused on I/EMRIs (Levin, 2007; Barausse et al., 2014; Derdzinski et al., 2019, 2021; Speri et al., 2023; Cole et al., 2023), recent works have indicated that measurements of gas effects on more equal-mass MBHBs are also possible (Garg et al., 2022; Dittmann et al., 2023; Tiede et al., 2023). However, these studies for MBHBs have focussed on the detectable accumulated dephasing in the GW waveform caused by gas, modelled using only Newtonian-order terms. This makes it optimistic as well as impossible to confidently pin down gas as a sole cause for this dephasing in the absence of an electromagnetic (EM) counterpart, since either higher-order post-Newtonian (PN) terms, eccentricity, or other environmental effects can also produce similar dephasings (see, e.g. Zwick et al., 2023). Therefore, it becomes crucial to identify the region of parameter space in which we could confidently identify a gas accretion disc as the environment of a MBHB, utilizing only the observed emitted GWs.
While shrinking the MBHB from a large scale, gas can also excite eccentricity that can be measurable up to $\sim 10^{-2.75}$ in the LISA band one year before the merger (Garg et al., 2024), despite partial circularization due to GWs (Peters & Mathews, 1963; Peters, 1964). Depending on whether the CBD is prograde or retrograde and extremely or moderately thin, and whether the binary is equal-mass or unequal-mass, we can expect different eccentricities in the LISA band. Therefore, measurement of eccentricities can provide evidence towards certain disc configurations even if gas effects themselves become negligible near coalescence. However, if there are measurable gas effects, then there can be an interplay between them and the eccentricity when performing parameter estimation.
@@ 2 The stationary phase approximation
Let us consider two spinning BHs in vacuum at redshift $z$, with a redshifted total binary mass $M_{z}$ and a primary-to-secondary mass ratio $q\geq 1$, revolving around each other in an eccentric orbit with their dimensionless spins [^1] The dimensionless spin of a BH of mass $m$ and spin angular momentum $J$ is $\chi\equiv cJ/Gm^{2}$, where $G$ is the gravitational constant and $c$ is the speed of light. $\chi_{1,2}$ aligned to the angular momentum of the binary. This is equivalent to the motion of a BH of reduced mass $\eta M_{z}$, where $\eta\equiv q/(1+q)^{2}$ is the symmetric mass ratio, in an elliptical orbit around a black hole of mass $M_{z}$, fixed at the focus. [^2] See Tables 1 and 7 for definitions of commonly used variables and terms. This orbit has a LISA frame semimajor axis $a$, eccentricity $e$, and orbital angular frequency $\Omega$.
@@
Since both eccentricity and environmental interactions affect the phase evolution of the source, we describe the cumulative phase of an event by its individual contributions: $\psi_{\rm TF2}$ is the phase a circular inspiral will accumulate in vacuum, solely due to emission of GWs; $\Delta\psi_{\rm TF2Ecc}$ represents the phase correction to an event’s waveform when orbital eccentricity alters its GW emission; and $\Delta\psi_{\rm gas}$ represents the phase correction that is a consequence of environmental interaction, which further speeds up or slows down the inspiral. For the latter, we adopt the interaction with a gas disk as our fiducial environmental effect. As we discuss later on, $\Delta\psi_{\rm TF2Ecc}$ and $\Delta\psi_{\rm gas}$ have negligible cross-terms. We consider terms up to $3.5$PN order (Buonanno et al., 2009) for $\psi_{\rm TF2}$ with aligned spin corrections also up to $3.5$PN order (Arun et al., 2009; Mishra et al., 2016), $3$PN and $\mathcal{O}(e^{2})$ order for $\Delta\psi_{\rm TF2Ecc}$ (Moore et al., 2016), and the leading-order correction from gas in $\Delta\psi_{\rm gas}$. The total SPA phase can be expressed as
@@
Given expected traditional disk model properties, we assume $\dot{a}_{\rm gas}\ll\dot{a}_{\rm GW}$ in the near-coalescence phase. Therefore, Eq. (4) can be expanded to separate the different contributions (all the integration constants are absorbed into $t_{c}$ and $\phi_{c}$):
@@ 3 Modeling effects from a circumbinary disc
The torque exerted by a near-equal mass MBHB typically carves out a central cavity in the inner accretion disc that leads to the formation of a circumbinary disc (CBD; D’Orazio et al. 2016). However, streams of gas still flow into this cavity, feeding mini-disks that accrete on to the binary as well as adding an additional torque component apart from an outer CBD (Farris et al., 2014). Non-axisymmetric features in this gas configuration exert a gravitational torque on the binary, which can lead to inspiral or outspiral of the MBHB depending on various disc-binary parameters. Typically a binary shrinks before GWs take over to drive it to merger, if we have a retrograde system (Tiede & D’Orazio, 2024), or a prograde disc with an unequal mass binary (Duffell et al., 2020), a sufficiently thin disc (Tiede et al., 2020), or a moderately eccentric system (D’Orazio & Duffell, 2021; Siwek et al., 2023). In 2D CBD studies, this gas torque is usually expressed as a function of the accretion rate ($\propto\dot{M}a^{2}\Omega$) and in the circular limit [^5] All quantities in the circular limit have bar on the top. as $\bar{\Gamma}_{\rm CBD}=\xi\dot{M}a^{2}\Omega$, where the fudge factor $\xi$ depends upon the disc parameters and binary mass ratio (Duffell et al., 2020; Garg et al., 2022). This expression is consistent with the viscous torque estimate (Lin & Papaloizou, 1986). Furthermore, $\xi$ tends to be positive (expand binary) for an equal-mass binary (D’Orazio & Duffell, 2021) and negative (shrink binary) for $q\gtrsim 10$ (Cuadra et al., 2009; Moody et al., 2019; Muñoz et al., 2019; Muñoz et al., 2020; Duffell et al., 2020; Dittmann et al., 2023; Tiede et al., 2020; Tiede & D’Orazio, 2024). However, depending on the thermodynamic assumptions, we can also have a negative torque for a circular near-equal mass system (Bourne et al., 2023). Therefore, in this work, we will survey $\xi$ values that cover all realistic possibilities.
@@ 3.1 Migration
For $\dot{a}_{\rm gas}=\dot{\bar{a}}_{\rm mig}$ in Eq. (8), we get the leading-order dephasing from migration by assuming a CBD torque fiducial model (see Appendix C):
@@
$C_{\rm mig}$ $=1.66\times 10^{-16}\frac{\xi}{1.0}\frac{{\rm f}_{\rm Edd}}{1.0}\frac{0.1}{\epsilon}\left(\frac{\eta}{0.25}\right)^{-2}\frac{M_{z}}{10^{5}{\rm M}_{\sun}},$
where $v\equiv(GM_{z}\pi f/c^{3})^{\frac{1}{3}}$ is the characteristic velocity of the binary, $\psi^{(0)}_{\rm TF2}\equiv(3/128\eta)v^{-5}$ is the leading-order phase term, ${\rm f}_{\rm Edd}$ is the Eddington ratio, [^8] We can express the accretion rate as $\dot{M}={\rm f}_{\rm Edd}\dot{M}_{\text{Edd}}$, where $\dot{M}_{\text{Edd}}=2.22\times 10^{-8}M(0.1/\epsilon){\rm yr}^{-1}$. and $\epsilon$ is the radiative efficiency, which we always set to $0.1$.
@@ 3.2 Mass accretion
The phase contribution due to mass accretion is (see Appendix C):
$\Delta\psi_{\rm acc}$ $=\psi^{(0)}_{\rm TF2}C_{\rm acc}v^{-8},$
$C_{\rm acc}$ $=-7.82\times 10^{-17}\frac{{\rm f}_{\rm Edd}}{1.0}\frac{0.1}{\epsilon}\left(\frac{\eta}{0.25}\right)^{-1}\frac{M_{z}}{10^{5}{\rm M}_{\sun}}.$

### v2

@@ Abstract
We assess the possibility of detecting both eccentricity and gas effects (migration and accretion) in the gravitational wave (GW) signal from LISA massive black hole binaries (MBHBs) at redshift $z=1$. Gas induces a phase correction to the GW signal with an effective amplitude ($C_{\rm g}$) and a semi-major axis dependence (assumed to follow a power-law with slope $n_{\rm g}$). We use a complete model of the LISA response, and employ a gas-corrected post-Newtonian in-spiral-only waveform model TaylorF2Ecc. By using the Fisher formalism and Bayesian inference, we constrain $C_{\rm g}$ together with the initial eccentricity $e_{0}$, the total redshifted mass $M_{z}$, the primary-to-secondary mass ratio $q$, the dimensionless spins $\chi_{1,2}$ of both component BHs, and the time of coalescence $t_{c}$. We find that simultaneously constraining $C_{\rm g}$ and $e_{0}$ leads to worse constraints on both parameters with respect to when considered individually. For a standard thin viscous accretion disc around $M_{z}=10^{5}~{}{\rm M}_{\sun}$, $q=8$, $\chi_{1,2}=0.9$, and $t_{c}=4$ years MBHB, we can confidently measure (with a relative error of $<50$ per cent) an Eddington ratio ${\rm f}_{\rm Edd}\sim 0.1$ for a circular binary and ${\rm f}_{\rm Edd}\sim 1$ for an eccentric system assuming $\mathcal{O}(10)$ stronger gas torque near-merger than at the currently explored much-wider binary separations. The minimum measurable eccentricity is $e_{0}\gtrsim 10^{-2.75}$ in vacuum and $e_{0}\gtrsim 10^{-2}$ in gas. A weak environmental perturbation (${\rm f}_{\rm Edd}\lesssim 1$) to a circular binary can be mimicked by an orbital eccentricity during in-spiral, implying that an electromagnetic counterpart would be required to confirm the presence of an accretion disc.
@@
pubyear: 2024pagerange: Measuring eccentricity and gas-induced perturbation from gravitational waves of LISA massive black hole binaries12
@@ 1 Introduction
The prospect of the observation of gravitational waves (GWs) in the mHz band in the 2030s looks promising following the adoption by ESA of the Laser Interferometer Space Antenna (LISA; Amaro-Seoane et al. 2017; Barack et al. 2019; Colpi et al. 2024) and with other projects, such as TianQin (Wang et al., 2019) and Taiji (Gong et al., 2021), being developed. One of the primary expected extragalactic sources for LISA are massive black hole binaries (MBHBs) with primary-to-secondary mass ratios $q\lesssim 10$ and total masses between $10^{4}~{}{\rm M}_{\sun}$ and $10^{8}~{}{\rm M}_{\sun}$, which LISA will be able to observe up to redshift $z\sim 20$ (Amaro-Seoane et al., 2017). Another expected source are intermediate/extreme mass ratio inspirals (I/EMRIs; Babak et al. 2017; Amaro-Seoane 2018) with $q\gtrsim 10^{2}$, which can be observed up to $z\lesssim 2$. MBHBs, with their high signal-to-noise ratios (SNRs; Amaro-Seoane et al. 2017), provide exciting opportunities to not only measure source properties with high accuracy but also place constraints on the properties of their environments.
@@
GWs can be an important tool to not only measure the source properties but also probe imprints of the environment in which the binary is evolving. In the coalescence phase, gas mainly affects the binary via migration torques and mass accretion. The detectability of the imprint on the emitted waveform of these effects (see, e.g. Barausse et al., 2014; Garg et al., 2022; Caputo et al., 2020) is dependent on the details of the gas inflow. While most of the works on environmental measurements from GWs have focused on I/EMRIs (Levin, 2007; Barausse et al., 2014; Derdzinski et al., 2019, 2021; Speri et al., 2023; Cole et al., 2023), recent works have indicated that measurements of gas effects on more equal-mass MBHBs are also possible (Garg et al., 2022; Dittmann et al., 2023; Tiede et al., 2024). However, these studies for MBHBs have focused on the detectable accumulated dephasing in the GW waveform caused by gas, modelled using only Newtonian-order terms. This makes it optimistic as well as impossible to confidently pin down gas as a sole cause for this dephasing in the absence of an electromagnetic (EM) counterpart, since either higher-order post-Newtonian (PN) terms, eccentricity, or other environmental effects can also produce similar dephasings (see, e.g. Zwick et al., 2023). Therefore, it becomes crucial to identify the region of parameter space in which we could confidently identify a gas accretion disc as the environment of a MBHB, utilizing only the observed emitted GWs.
While shrinking the MBHB from a large scale, gas can also excite eccentricity that can be measurable up to $\sim 10^{-2.75}$ in the LISA band one year before the merger (Garg et al., 2024b), despite partial circularization due to GWs (Peters & Mathews, 1963; Peters, 1964). Depending on whether the CBD is prograde or retrograde and extremely or moderately thin, and whether the binary is equal-mass or unequal-mass, we can expect different eccentricities in the LISA band. Therefore, measurement of eccentricities can provide evidence towards certain disc configurations even if gas effects themselves become negligible near coalescence. However, if there are measurable gas effects, then there can be an interplay between them and the eccentricity when performing parameter estimation.
@@ 2 The stationary phase approximation
Let us consider two spinning BHs in vacuum at redshift $z$, with a redshifted total binary mass $M_{z}$ and a primary-to-secondary mass ratio $q\geq 1$, revolving around each other in an eccentric orbit with their dimensionless spins [^1] The dimensionless spin of a BH of mass $m$ and spin angular momentum $J$ is $\chi\equiv cJ/Gm^{2}$, where $G$ is the gravitational constant and $c$ is the speed of light. $\chi_{1,2}$ aligned to the angular momentum of the binary. This is equivalent to the motion of a BH of reduced mass $\eta M_{z}$, where $\eta\equiv q/(1+q)^{2}$ is the symmetric mass ratio, in an elliptical orbit around a black hole of mass $M_{z}$, fixed at the focus. [^2] See Tables 1 and 7 for definitions of commonly used variables and terms. This orbit has a detector frame semimajor axis $a$, eccentricity $e$, and orbital angular frequency $\Omega$.
@@
Since both eccentricity and environmental interactions affect the phase evolution of the source, we describe the cumulative phase of an event by its individual contributions: $\psi_{\rm TF2}$ is the phase a circular inspiral will accumulate in vacuum, solely due to emission of GWs; $\Delta\psi_{\rm TF2Ecc}$ represents the phase correction to an event’s waveform when orbital eccentricity alters its GW emission; and $\Delta\psi_{\rm gas}$ represents the phase correction that is a consequence of environmental interaction, which further speeds up or slows down the inspiral. For the latter, we adopt the interaction with a gas disc as our fiducial environmental effect. As we discuss later on, $\Delta\psi_{\rm TF2Ecc}$ and $\Delta\psi_{\rm gas}$ have negligible cross-terms. We consider terms up to $3.5$PN order (Buonanno et al., 2009) for $\psi_{\rm TF2}$ with aligned spin corrections also up to $3.5$PN order (Arun et al., 2009; Mishra et al., 2016), $3$PN and $\mathcal{O}(e^{2})$ order for $\Delta\psi_{\rm TF2Ecc}$ (Moore et al., 2016), and the leading-order correction from gas in $\Delta\psi_{\rm gas}$. The total SPA phase can be expressed as
@@
Given expected traditional disc model properties, we assume $\dot{a}_{\rm gas}\ll\dot{a}_{\rm GW}$ in the near-coalescence phase. Therefore, Eq. (4) can be expanded to separate the different contributions (all the integration constants are absorbed into $t_{c}$ and $\phi_{c}$):
@@ 3 Modeling effects from a circumbinary disc
The torque exerted by a near-equal mass MBHB typically carves out a central cavity in the inner accretion disc that leads to the formation of a circumbinary disc (CBD; D’Orazio et al. 2016). However, streams of gas still flow into this cavity, feeding mini-discs that accrete on to the binary as well as adding an additional torque component apart from an outer CBD (Farris et al., 2014). Non-axisymmetric features in this gas configuration exert a gravitational torque on the binary, which can lead to inspiral or outspiral of the MBHB depending on various disc-binary parameters. Typically a binary shrinks before GWs take over to drive it to merger, if we have a retrograde system (Tiede & D’Orazio, 2024), or a prograde disc with an unequal mass binary (Duffell et al., 2020), a sufficiently thin disc (Tiede et al., 2020), or a moderately eccentric system (D’Orazio & Duffell, 2021; Siwek et al., 2023). In 2D CBD studies, this gas torque is usually expressed as a function of the accretion rate onto the binary [^5] We can express $\dot{M}_{z}={\rm f}_{\rm Edd}M_{z}/\tau$, where ${\rm f}_{\rm Edd}$ is the Eddington ratio, and $\tau\approx 50$Myr is the Salpeter timescale for our fiducial radiative efficiency of $0.1$. ($\dot{M}_{z}$) and in the circular limit [^6] All quantities in the circular limit have bar on the top. as $\bar{\Gamma}_{\rm CBD}=\xi\dot{M}_{z}a^{2}\Omega$, where the fudge factor $\xi$ depends upon the disc parameters and binary mass ratio (Duffell et al., 2020; Garg et al., 2022). This expression is consistent with the viscous torque estimate (Lin & Papaloizou, 1986). Furthermore, $\xi$ tends to be positive (expand binary) for an equal-mass binary (D’Orazio & Duffell, 2021) and negative (shrink binary) for $q\gtrsim 10$ (Cuadra et al., 2009; Moody et al., 2019; Muñoz et al., 2019; Muñoz et al., 2020; Duffell et al., 2020; Dittmann et al., 2023; Tiede et al., 2020; Tiede & D’Orazio, 2024). However, depending on the thermodynamic assumptions, we can also have a negative torque for a circular near-equal mass system (Bourne et al., 2023). Therefore, in this work, we will survey $\xi$ values that cover all realistic possibilities.
@@ 3.1 Migration
For $\dot{a}_{\rm gas}=\dot{\bar{a}}_{\rm mig}$ and assuming $\dot{a}_{\rm GW}=\dot{a}^{(0)}_{\rm GW}$ in Eq. (8), we get the leading-order dephasing from migration (Yunes et al., 2011)
$\Delta\psi_{\rm mig}$ $=-\psi^{(0)}_{\rm TF2}\frac{{20\mathcal{A}}}{(n_{\rm g}+4)(2n_{\rm g}+5)}v^{-2n_{\rm g}}.$ (10)
Assuming a thin CBD torque fiducial model, the torque on the binary in the circular limit is given by:
$\bar{\Gamma}_{\rm CBD}$ $=\xi\dot{M}a^{2}\Omega=\frac{d}{dt}(\eta M_{z}a^{2}\Omega),$
$\implies\dot{\bar{a}}_{\rm mig}$ $=\xi\frac{\dot{M}}{M_{z}}\frac{2}{\eta}a,$ (11)
where in the first line we have ignored the mass accretion term when taking the time-derivative (see Section 3.2). Therefore, equating Eqs (9) and (3.1) gives us
${\mathcal{A}}$ $=-5.40\times 10^{-15}\frac{\xi}{1.0}\frac{{\rm f}_{\rm Edd}}{1.0}\frac{0.1}{\epsilon}\left(\frac{\eta}{0.1}\right)^{-2}\frac{M_{z}}{10^{5}{\rm M}_{\sun}},$ (12)
$n_{\rm g}$ $=4.$
Plugging $\mathcal{A}$ and $n_{\rm g}$ into Eq. (10) gives us
@@
$C_{\rm mig}$ $=1.04\times 10^{-15}\frac{\xi}{1.0}\frac{{\rm f}_{\rm Edd}}{1.0}\frac{0.1}{\epsilon}\left(\frac{\eta}{0.1}\right)^{-2}\frac{M_{z}}{10^{5}{\rm M}_{\sun}},$
where $v\equiv(GM_{z}\pi f/c^{3})^{\frac{1}{3}}$ is the characteristic velocity of the binary, $\psi^{(0)}_{\rm TF2}\equiv(3/128\eta)v^{-5}$ is the leading-order phase term.
@@ 3.2 Mass accretion
The accretion rate onto the binary for $\epsilon=0.1$ can be expressed as
$\dot{M}_{z}=$ ${\rm f}_{\rm Edd}M_{z}/\tau,$
$\implies M_{z}=$ $M_{z,0}\exp({\rm f}_{\rm Edd}t/\tau).$ (14)
The phase contribution due to mass accretion in the circular limit can be expressed as (Caputo et al., 2020):
$\Delta\psi_{\rm acc}=\psi^{(0)}_{\rm TF2}\frac{25}{256}\frac{{\rm f}_{\rm Edd}}{\tau}\frac{GM_{z}\eta^{-1}}{c^{3}}\left(\frac{1}{3}v_{0}^{-8}-\frac{15}{26}v^{-8}\right),$ (15)
where we have absorbed terms independent of $f$ into $\phi_{c}$ and terms proportional to $f$ in $t_{c}$. We have replaced $M_{z,0}$ with $M_{z}$ after computations, since the LISA observation duration we consider is at maximum $4~{}{\rm years}$, implying ${\rm f}_{\rm Edd}t/\tau\ll 1$ in Eq. (3.2). First term in $\Delta\psi_{\rm acc}$ can be re-expressed as
$\Delta\psi^{(I)}_{\rm acc}$ $\approx\psi^{(0)}_{\rm TF2}\times 10^{-16}\frac{{\rm f}_{\rm Edd}}{\eta}\frac{M_{z}}{10^{5}{\rm M}_{\sun}}\left(\frac{a_{0}}{GM/c^{2}}\right)^{4},$
where $a_{0}$ is the initial separation, which cannot be more than $\sim\mathcal{O}(100GM_{z}/c^{2})$ to have the MBHB merge within the LISA observation window. Therefore, $\Delta\psi^{(I)}_{\rm acc}$ is at maximum $\mathcal{O}(10^{-6})\psi^{(0)}_{\rm TF2}$ even for optimistic accretion rates (${\rm f}_{\rm Edd}$ $\lesssim 100$) for near equal-mass MBHBs. Hence, the first term in Eq. (15) can be dropped due to being negligible with respect to the leading-order SPA phase contribution $\psi^{(0)}_{\rm TF2}$ and we get the effective accretion dephasing
$\Delta\psi_{\rm acc}$ $=\psi^{(0)}_{\rm TF2}C_{\rm acc}v^{-8},$ (16)
$C_{\rm acc}$ $=-1.96\times 10^{-16}\frac{{\rm f}_{\rm Edd}}{1.0}\frac{0.1}{\epsilon}\left(\frac{\eta}{0.1}\right)^{-1}\frac{M_{z}}{10^{5}{\rm M}_{\sun}}.$